### v1

# AgileFormer: Spatially Agile Transformer UNet for Medical Image Segmentation

Abstract
In the past decades, deep neural networks, particularly convolutional neural networks, have achieved state-of-the-art performance in a variety of medical image segmentation tasks. Recently, the introduction of the vision transformer (ViT) has significantly altered the landscape of deep segmentation models. There has been a growing focus on ViTs, driven by their excellent performance and scalability. However, we argue that the current design of the vision transformer-based UNet (ViT-UNet) segmentation models may not effectively handle the heterogeneous appearance (e.g., varying shapes and sizes) of objects of interest in medical image segmentation tasks. To tackle this challenge, we present a structured approach to introduce spatially dynamic components to the ViT-UNet. This adaptation enables the model to effectively capture features of target objects with diverse appearances. This is achieved by three main components: (i) deformable patch embedding; (ii) spatially dynamic multi-head attention; (iii) deformable positional encoding. These components were integrated into a novel architecture, termed AgileFormer. AgileFormer is a spatially agile ViT-UNet designed for medical image segmentation. Experiments in three segmentation tasks using publicly available datasets demonstrated the effectiveness of the proposed method. The code is available at https://github.com/sotiraslab/AgileFormer.
Keywords Transformer $\cdot$ UNet $\cdot$ Medical Image Segmentation.
Figure 1: (a) Comparison of model scalability in the Synapse dataset. The base model is almost four times larger than the tiny model. (b) A roadmap going from SwinUNet to the proposed AgileFormer: from the top to the bottom, each row represents a model design variant. The foreground bars are DSC in the FLOP regime of different design variants; a hatched bar means the modification leads to a performance drop.
1 Introduction
Medical image segmentation tasks are important in modern medicine, as it is typically the first step in many image-driven diagnoses and analyses [1, 2]. Deep learning based automated segmentation methods have dominated this field due to their high efficiency and state-of-the-art performance. Among them, convolutional neural networks (CNNs) [3, 4, 5, 6, 7, 8] have emerged as the most prevalent choice since the introduction of UNet [3]. This is due to CNNs’ inherent advantages in handling image-driven tasks, such as their ability to capture locality and translation invariance. However, they struggle with capturing global semantics primarily due to their restricted receptive field.
In contrast, the recently proposed vision transformer (ViT) [9] mitigates this problem through a self-attention mechanism that captures dependencies between image patches regardless of their spatial distances. The first ViT-based UNet (ViT-UNet) for medical image segmentation combined a ViT encoder with a CNN decoder, and hence was termed TransUNet [1]. However, TransUNet is buderened by a large number of parameters ($\sim$100M) and considerable computational complexity. This is because it employs standard self-attention, which has a quadratic time and memory complexity w.r.t. the input token size. Leveraging the window attention [10], which performs self-attention within a small window in parallel for all image patches, SwinUNet [2] mitigates this challenge and is the first pure ViT-UNet, with self-attention as the main feature extractor. However, SwinUNet uses fixed-size windowing, which may limit its ability to capture precise representation for target objects with varying sizes and shapes. This potentially limits its adaptability and generalizability to multi-class segmentation tasks.
Additionally, SwinUNet do not exhibit good scaling behavior. In other words, their performance did not scale effectively when increasing the size of the model [2] (as evident in Figure˜1(a)). This is in contrast to evidence provided by empirical studies using natural images that have revealed that ViTs exhibit exceptional scaling behavior. We hypothesize this phenomenon may also be attributed to heterogeneity in the size and shape of objects targeted for segmentation, which cannot be handled by fixed-size windowing. To address these challenges, spatially dynamic components are necessary for improving the performance and scalability of a ViT-UNet. Although recent works have explored introducing dynamic (e.g., multi-scale [11], deformable [12, 13]) components into ViT-UNets for medical imaging, they have mainly focused on adopting existing deformable attention. However, they did not investigate introducing dynamic components into other parts (i.e., patch embedding and positional encoding) of a ViT-UNet. We argue that enhancing these other components with dynamic elements is equally important. We kindly direct the readers to Appendix˜A for an in-depth discussion on how AgileFormer fundamentally differs from the three aforementioned works  [11, 12, 13].
Accordingly, we developed a novel spatially agile pure ViT-UNet to capture diverse target objects in a medical image segmentation task. First, we replace the standard rigid square patch embedding in ViT-UNet with a novel deformable patch embedding. Second, we adopt spatially dynamic self-attention [14] as the building block to capture spatially varying features. Third, we propose a novel multi-scale deformable positional encoding to model irregularly sampled grids in self-attention. We integrated these dynamic components into a novel ViT-UNet architecture, named AgileFormer. Extensive experiments demonstrated the effectiveness of the proposed method for three medical image segmentation tasks. AgileFormer outperformed recent state-of-the-art UNet models for medical image segmentation and demonstrated exceptional scalability.
2 Method
In this section, we provide a roadmap going from the standard SwinUNet to the proposed AgileFormer (see Figure˜1(b)). To make this paper self-contained, we first provide a brief introduction of the essential elements in ViT-UNets. A ViT-UNet (e.g., SwinUNet) is a U-shaped encoder-decoder network with skip connections, wherein the primary feature extractions in both the encoder and decoder are achieved by the self-attention mechanism. The building block of a standard ViT-UNet is composed of three fundamental components: patch embedding, self-attention, and positional encoding. The patch embedding projects image patches into feature embeddings. More recent ViTs [15, 14] even treat downsampling (i.e., patch merging in a SwinUNet) as part of the patch embedding. In this paper, we will follow the same convention. The self-attention, which captures the dependencies between image patches, is used for the main feature extraction. For computational tractability and the need of locality, recent ViT-UNets use window-based self-attention mechanism [2, 16, 13, 17]. Unlike convolution, self-attention discards the spatial correlation between image patches, which hinders the localization ability of pure ViT in segmentation task. The positional encoding is used to address this limitation. We would like to point out that there are also hybrid models [1, 16, 18, 17, 12, 19], where the main feature extraction in the encoder/decoder is achieved by both CNN and ViT. Whereas, we focus on developing a pure ViT-UNet in this paper.
Figure 2: The architecture of the proposed AgileFormer.
The rest of this section is organized to introduce the three main components of the proposed AgileFormer in Figure˜2: deformable patch embedding (Section˜2.1), spatially dynamic multi-head attention (Section˜2.2), and deformable positional encoding (Section˜2.3).
2.1 Deformable Patch Embedding
2.1.1 Rigid patch embedding
The ViT-UNet starts by converting image patches into tokens. This process typically involves partitioning an image into a sequence of non-overlapping $n\times n\ (\times n)$ (e.g., $4\times 4\ (\times 4)$ in SwinUNet [2, 20, 17]) fixed-size patches. Subsequently, each of these patches is projected into a 1D feature vector. The main reason for performing a rigid (square) patch embedding is due to its simplicity, as it can be easily implemented as a standard convolution (kernel size $k=n$; stride $s=n$; dilation $d=1$). However, we argue that this rigid patch embedding is not optimal for a segmentation task for two main reasons. First, segmentation requires more precise pixel-level localization. However, the rigid patch embedding can only provide patch-level localization. Second, the shape and size of target objects in most medical image segmentation tasks (e.g., the multi-organ segmentation task) vary significantly.
2.1.2 Deformable patch embedding
To address the limitation of rigid patch embedding, we propose a deformable patch embedding (see Figure˜2) by leveraging the deformable convolution [21]. The deformable convolution is defined as
$(\boldsymbol{f}\ast\boldsymbol{k})[p]=\sum_{p_{k}\in\Omega}\phi(\boldsymbol{f};p+\Delta p+p_{k})\cdot\boldsymbol{k}[p_{k}],$ (1)
where $\boldsymbol{f}\in\mathbb{R}^{L\times d_{f}}$ is a $d_{f}$-dimensional feature map with a uniform grid of $L$ locations $p\in\mathbb{R}^{L\times D}$ ($D=2$ for 2D; $D=3$ for 3D). $\boldsymbol{k}$ is the convolutional kernel that operates on the grid $\Omega=[p_{k}]$ defines $k$-nearest (i.e., kernel size) neighboring locations of $p$. $\Delta p]$ are offsets from which the irregular grid is sampled. The offsets are learned through a single convolutional layer $\Delta p=\operatorname{Conv}_{\text{offset}}(\boldsymbol{f})$. $\phi$ is a sampling function that performs bilinear/trilinear interpolation to sample locations $p+\Delta p+p_{k}]$ in $\boldsymbol{f}$, as the offsets $\Delta p$ are typically fractional.
First patch embedding. We replaced the single-layer rigid convolution patch embedding in a standard SwinUNet with two consecutive deformable convolutional layers ($k=3$; $s=n/2$; $d=1$). The rationale behind this is that the two consecutive overlapping deformable patch embeddings can extract better local representations, which compensates for the lack of locality in the self-attention.
Downsampling patch embedding. We also replaced the patch merging in standard SwinUNet with convolutional downsampling via a single layer ($k=3$; $s=2$; $d=1$) for downsampling. We used overlapping kernels to better preserve localized patterns [22, 15, 20], which also aligns with the overlapping deformable patch embedding.
2.2 Spatially Dynamic Self-Attention
The self-attention is the building block of a ViT UNet. Unlike convolution, self-attention does not enforce any spatial inductive bias but makes decisions purely by relying on dependencies (mainly similarities) between tokens while lacking the ability to capture spatially adaptive features for multi-class segmentation. Accordingly, we propose using a spatially dynamic self-attention module as the building block of ViT-UNet. This module is inspired from [15], and it includes deformable multi-head attention (DMSA) [15] and neighborhood multi-head attention (NMSA) [23]. The transformer block is constructed by alternating these two attention mechanisms (see Figure˜2). We also distribute more computation to the third encoder block with a stage ratio of $[1:2:5:1]$, instead of $[2:2:2:2]$ in SwinUNet. This is because the third layer of the encoder typically captures better feature representations than the others [24].
2.2.1 Deformable Multi-head Self-Attention (DMSA)
The deformable multi-head (i.e., $H$ head) attention [15] for the $h$-th head is formulated as
$\texttt{DMSA}_{h}(\boldsymbol{f})=\operatorname{softmax}(\boldsymbol{Q}_{h}\tilde{\boldsymbol{K}}_{h}^{\top}/\sqrt{d_{k}})\tilde{\boldsymbol{V}}_{h}$ (2)
where:
$\boldsymbol{Q_{h}}=\boldsymbol{f}\boldsymbol{W}_{h}^{Q},\quad\tilde{\boldsymbol{K}}_{h}=\tilde{\boldsymbol{f}}\boldsymbol{W}_{h}^{K},\quad\tilde{\boldsymbol{V}}_{h}=\tilde{\boldsymbol{f}}\boldsymbol{W}_{h}^{V},\quad\tilde{\boldsymbol{f}}=\phi(\boldsymbol{f};p+\Delta p_{h}).$ (3)
Here, we reuse the notation in Equation˜1 with $p$ being a uniform grid of points, $\Delta p_{h}$ being the generated offsets for the $h$-th head, and $\phi$ being an interpolation function. $\{\boldsymbol{W}_{h}^{Q},\boldsymbol{W}_{h}^{K}\}\in\mathbb{R}^{d_{f}\times d_{k}}$ and $\boldsymbol{W}_{h}^{V}\in\mathbb{R}^{d_{f}\times d_{v}}$ are trainable parameters, and $d_{k}$, $d_{v}$ are the hidden dimensions of linear projection of the key $\tilde{\boldsymbol{K}}_{h}$ and the value $\tilde{\boldsymbol{V}}_{h}$ in DMSA. The offsets in DMSA are also generated by passing the query through a convolutional layer $\Delta p=\operatorname{Conv}_{\text{offset}}(\boldsymbol{Q}_{h})$. The resulting irregularly sampled feature map is denoted as $\tilde{\boldsymbol{f}}$ in Equation˜3. Similar to deformable convolution, the irregularly sampled feature map is then applied to self-attention by using irregularly sampled key $\tilde{\boldsymbol{K}}_{h}$ and value $\tilde{\boldsymbol{V}}_{h}$ (see Equation˜2).
2.2.2 Neighborhood Multi-head Self-Attention (NMSA)
In contrast to standard self-attention, which computes the similarity of each element at a given position $p$ on a feature map $\boldsymbol{f}$ with every other element, the construction of neighborhood attention [23] only leverages the information from the $k$-nearest neighbors around location $p$. We reuse the notation $k$ in Equation˜2, as the NMSA operates like a convolution. As a result, neighborhood attention reduces the quadratic computational complexity in standard self-attention to approximately linear w.r.t. the spatial dimension of $\boldsymbol{f}$, as $k$ is typically small (e.g., $k=7\times 7\ (\times 7)$). Furthermore, this reintroduces local operations into self-attention, allowing for translational equivariance, and thereby enhancing the ability to better preserve local information. Following the notation in Equation˜2, the neighborhood multi-head attention at position $p_{l}$ is computed as
[…]
where:
$\begin{split}\hat{\boldsymbol{K}}_{h}[p_{l}]&=[\boldsymbol{K}_{h}[p_{l1}],\cdots,\boldsymbol{K}_{h}[p_{lk}]],\quad\boldsymbol{K}_{h}=\boldsymbol{f}\boldsymbol{W}_{h}^{K}\\ \hat{\boldsymbol{V}}_{h}[p_{l}]&=[\boldsymbol{V}_{h}[p_{l1}]^{\top},\cdots,\boldsymbol{V}_{h}[p_{lk}]^{\top}]^{\top},\quad\boldsymbol{V}_{h}=\boldsymbol{f}\boldsymbol{W}_{h}^{V}.\end{split}$ (5)
Here, $p_{lk}]$ denotes the $k$-th neighboring locations for a given location $p_{l}$. It is worth noting that the dimension of the resulting attention is $\mathbb{R}^{L\times K}$ with $K=k\times k(\times k)$, instead of $\mathbb{R}^{L\times L}$ in standard self-attention or window-attention.
2.3 Multi-scale Deformable Positional Encoding
Figure 3: The proposed multi-scale deformable positional encoding for irregularly sampled grids in deformable multi-head self-attention.
The design of positional encoding (PE) is barely explored by previous ViT UNets. Most ViT-UNets either disregard positional encoding [1] or inherit PEs from their ancestor models [25, 2, 16]. Specifically, [25] used an absolute PE (APE), which assigns an absolute value for each token. While the others [2, 16, 20] used a relative PE (RPE) [26] to encode the relative positions between tokens. However, these are designed for 1D signals while neglecting the spatial correlation. Consequently, they are not well adapted for modeling 2D/3D signals with spatial correlations. Recently, conditional PE (CPE) [27] was designed for vision tasks offering the ability of both APE and RPE at the same time. More importantly, standard APE, RPE, and CPE are not directly applicable to irregularly sampled grids, as all of them are formulated in a rigid grid. In accordance with the irregularly sampled DMSA presented in this paper, we proposed a multi-scale deformable positional encoding (MS-DePE) designed to encode irregularly sampled positional information across various scales (see Figure˜3).
The proposed MS-DePE is formulated in a conditional positional encoding fashion [27] as follows:
$\texttt{MS-DePE}(\boldsymbol{f})=\boldsymbol{f}+\texttt{P}_{\theta}(\boldsymbol{f}),$ (6)
where $\boldsymbol{f}$ is the input feature map, and $\texttt{P}_{\theta}$ denotes the trainable positional encoding layer parameterized by $\theta$. $\texttt{P}_{\theta}$ is implemented as multi-scale deformable depth-wise convolutional layers with different kernel sizes (i.e., $3\times 3\ (\times 3)$ and $5\times 5\ (\times 5)$). For this purpose, we first recover the spatial resolution of the feature map $\boldsymbol{f}$ before passing it through $\texttt{P}_{\theta}$. After applying $\texttt{P}_{\theta}$, we flatten the feature map back to its original shape (see Fig. 3 for details).
2.4 Model construction
Similar to [1, 2, 11], we developed two variants of AgileFormer by varying the embedding dimension ($d_{f}$) and the number of heads ($H$), but kept the main structures unchanged (i.e., number of transformer blocks in the encoder/decoder): AgileFormer-T(iny) ($d_{f}=[64,128,256,512]$; $H=[2,4,8,16]$) and AgileFormer-B(ase) ($d_{f}=[128,256,512,1024]$; $H=[4,8,16,32]$). We also incorporated deep supervision (DS) outlined in [7, 20] into the proposed method.
3 Experiments and results
3.1 Experimental design
Dataset. In line with previous works [1, 2, 20], we validated the effectiveness of the proposed method on three publicly available medical image segmentation datasets: the Synapse dataset [28], the Automated Cardiac Diagnosis Challenge (ACDC) dataset [29], and the brain tumor segmentation dataset from the Decathlon challenge [30]. The Synapse multi-organ dataset includes 30 3D abdominal CT scans with corresponding segmentation masks of 8 organs (i.e., aorta, gallbladder, kidney left, kidney right, liver, pancreas, spleen, stomach). The ACDC dataset consists of 100 3D cardiac MRI scans, each with a segmentation mask that features the right ventricle (RV), myocardium (Myo), and left ventricle (LV). The Decathlon brain tumor dataset comprises 484 3D multi-modal brain tumor MRI scans with segmentation masks delineating enhancing tumor, non-enhancing tumor, and edema. To be consistent with previous works [18, 20], we reported the results of the whole tumor (WT), enhancing tumor (ET) and tumor core (TC).
Experimental setup and evaluation metric. We adhered to the experimental protocols outlined in [1, 2] for the Synapse and ACDC datasets, including training/testing sets, input image size (i.e., $224\times 224$), data augmentations, model selection, and evaluations. For a fair comparison with the previous literature, experiments for Synapse and ACDC datasets were carried out in 2D. For the 3D volumetric brain tumor segmentation, we followed the experimental protocols specified in [18] for a fair comparison with the previous literature. The input image was set to $96\times 96\times 96$. We evaluated the segmentation performance using the dice similarity coefficient (DSC). Statistical significance between the average DSC of the best and second-best model was estimated using the paired t-test. We emphasize that all competing methods applied in the Synapse, ACDC, and Decathlon datasets followed the same experimental protocols.
Table 1: Performance comparison with 2D methods in the multi-organ segmentation using the Synapse dataset. The best result within each column is highlighted by bold, and the second-best is highlighted with an underline. : models implemented by us. The other benchmarks were taken from [1, 2, 11, 19, 16, 31, 32, 33]. (: $p<0.01$; with the paired t-test to the second-best model. : Large models that have a similar parameters to the AgileFormer-B model (TransCASCADE: 123.47M; MERIT: 147.86M; AgileFormer-B: 112.97M), where deep supervision is also applied.)
[…]
CNN
U-Net [3] 76.85 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
R50 Att-UNet [1] 75.57 55.92 63.91 79.20 72.71 93.56 49.37 87.19 74.95
Att-UNet [4] 77.77 89.55 68.88 77.98 71.11 93.57 58.04 87.30 75.75
Hybrid
TransUNet [1] 77.48 87.23 63.13 81.87 77.02 94.08 55.86 85.08 75.62
MixedUNet [19] 78.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
CoTr [12] 78.56 87.09 65.37 86.19 80.32 94.22 52.28 87.01 76.00
Hiformer [16] 80.29 85.63 73.29 82.39 64.84 94.22 60.84 91.03 78.07
TransCASCADE [33] 82.68 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.69
ViT
R50 ViT [1] 71.29 73.73 55.13 75.80 72.20 91.51 45.99 81.99 73.95
SwinUNet [2] 79.13 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
TransDeepLab [34] 80.16 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40
SDAUT [13] 80.67 87.03 69.30 81.87 80.20 94.91 64.56 89.54 77.91
CATFormer [35] 82.17 88.98 67.16 85.72 81.69 95.34 66.53 90.74 81.20
MissFormer [31] 81.96 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
MERIT [11] 84.90 87.71 74.40 87.79 84.85 95.26 71.81 92.01 85.38
MERIT 82.89 87.42 72.61 87.94 85.36 94.45 67.63 87.15 80.53
AgileFormer-T w/o DS 82.74 88.08 75.16 82.41 81.36 95.09 67.23 90.94 81.61
AgileFormer-T w/ DS 83.59 88.81 74.43 84.61 82.78 95.48 69.45 90.14 83.05
AgileFormer-B w/o DS 84.14 87.76 74.71 86.69 83.81 95.31 68.28 91.00 85.58
AgileFormer-B w/ DS 85.74 89.11 77.89 88.83 85.00 95.64 71.62 92.20 85.63
AgileFormer-B w/ DS [best] 86.11 89.29 78.51 89.09 85.12 95.96 71.52 93.03 86.35
Note: MERIT uses the testing set to select the best model, which is not fair for comparison. The evaluation of MERIT using the last-epoch checkpoint results in a lower performance. AgileFormer can even achieve higher performance when using the testing set for model selection. We report this result as AgileFormer-B w/s DS [best] only for comparison, but we highly disencourage any reader from citing this result.
Table 2: Performance comparison with 2D methods in cardiac MRI segmentation using the ACDC dataset. : models implemented by us. The other benchmarks were taken from [1, 2, 19, 31, 33, 11].
[…]
TransUNet [1] 89.71 88.86 84.53 95.73
CoTr [12] 90.52 87.81 88.44 95.29
SwinUNet [2] 90.00 88.55 85.62 95.83
MixedUNet [19] 90.43 86.64 89.04 95.62
MissFormer [31] 90.86 89.55 88.04 94.99
SDAUT [13] 91.08 89.37 88.58 95.28
PVT-CASCADDE [33] 91.46 88.00 89.97 95.50
TransCASCADE [33] 91.63 89.14 90.25 95.50
MERIT [11] 92.32 90.87 90.00 96.08
AgileFormer-T w/ DS 91.76 89.80 89.71 95.77
AgileFormer-B w/ DS 92.55 91.05 90.40 96.19
Table 3: Performance comparison with 3D models in brain tumor segmentation using the Decathlon dataset. : models implemented by us. The input size was set to $96\times 96\times 96$. Please refer to Appendix˜A for detials of implementing the 3D experiments. We only compared the tiny model (AgileFormer-T), as it has a similar scale to the baselines.
[…]
TransUNet [1] 84.6 90.5 79.9 83.4
TransBTS [25] 84.2 90.2 79.6 82.9
CoTr [12] 84.1 91.2 78.7 83.2
UNETR [18] 84.1 90.5 79.3 82.4
SwinUNETR [17] 84.9 90.7 80.5 83.5
nnFormer [20] 84.9 91.4 79.2 84.0
AgileFormer-T w/ DS 85.7 91.2 80.8 85.1
Implementation details. All models were trained using a combination of dice and cross-entropy loss [2, 20], employing an AdamW [36] optimizer with cosine learning rate decay and a 20-epoch warmup. We initialized model parameters with ImageNet pre-trained weights. In the 2D experiments, a batch size of 24 was used, the initial learning rate was set to $3\times 10^{-4}$, and the model was trained for 400 epochs. For the 3D experiments, the batch size and learning rate were set to 4 and $5\times 10^{-4}$, and the model was trained for 200 epochs. We would like to point out that we reimplemented all the 3D baseline models using the nnFormer framework (see Appendix˜B for details of 3D experiments). All experiments were implemented using PyTorch and were performed on a Nvidia V100 GPU with 32GB memory.
Figure 4: Visual comparison of (a) multi-organ, (b) cardiac, and (c) brain tumor segmentation across Synapse, ACDC, and Decathlon datasets. AgileFormer shows the best qualitative segmentation quality across all three datasets.
3.2 Main results
Our proposed AgileFormer outperformed recent state-of-the-art (SOTA) methods in all three datasets (Table 13 and 3). Specifically, the proposed method performed better than the second-best model by 0.84, 0.23, and 0.80 in average DSC (%) for Synapse multi-organ, ACDC cardiac, and brain tumor segmentation tasks, respectively. This improvement was shown to be statistically significant. Specifically, the proposed method achieved an average DSC of 85.74, surpassing the other baseline methods in segmenting all 6 out of 8 organs in the Synapse dataset (Table˜1). For the remaining two organs, the proposed method is the second-best performing method. The proposed method achieved an average DSC of 92.55 in the ACDC dataset, surpassing all the baseline methods in segmenting RV, Myo, and LV (see Table˜3). In the 3D volumetric brain tumor segmentation task, the proposed method achieved an average DSC of 85.7, surpassing recent SOTA ViT-UNet methods (Table˜3). In addition, the proposed method demonstrated the best qualitative segmentation quality across these three public datasets (see  Figure˜4).
3.3 Ablations on model design variants
Table 4: Ablation studies on model design variants in Synapse multi-organ segmentation task, starting from a SwinUNet baseline. This is also the roadmap going from SwinUNet to the proposed AgileFormer.
[…]
Baseline (WMSA + RPE) 79.13 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60 27.17 / 6.13
Convolutional Downsampling 79.40 85.74 70.56 82.33 80.69 93.50 58.89 91.46 72.05 29.10 / 6.34
Deformable Patch Embedding 80.91 87.38 71.25 85.17 83.08 94.11 58.71 88.92 78.67 29.14 / 6.54
NMSA + WMSA (w/o PE) 81.27 87.22 68.79 84.73 79.49 95.31 63.13 90.62 80.86 26.74 / 6.31
WMSA + DMSA (w/o PE) 81.47 86.12 67.37 84.18 80.13 95.61 64.86 90.97 82.55 26.77 / 5.98
NMSA + DMSA (w/o PE) 82.06 87.54 70.95 81.90 79.29 95.36 68.03 90.72 82.72 26.82 / 5.99
+ APE 81.96 88.35 70.72 81.82 78.84 95.09 67.80 90.27 82.45 28.90 / 5.99
APE $\rightarrow$ deformable RPE [15] 82.42 87.17 73.32 86.08 82.54 95.08 65.65 88.64 80.92 26.99 / 6.00
deformable RPE $\rightarrow$ CPE 81.91 88.13 72.28 79.81 77.75 95.06 65.24 91.90 85.14 26.87 / 6.01
CPE $\rightarrow$ MS-DePE 82.74 88.08 75.16 82.41 81.36 95.09 67.23 90.94 81.61 27.47 / 7.08
+ Deep Supervision 83.59 88.81 74.43 84.61 82.78 95.48 69.45 90.14 83.05 28.85 / 7.36
We conducted ablation studies on the model design variants, including patch embedding and different choices of spatially dynamic attention and positional encoding (see Table˜4). First, we showed that replacing patch merging with convolutional downsampling led to a small performance gain of $0.34\%$ in average DSC. Then, adding deformable embedding can lead to a performance gain of $1.9\%$ in DSC. Second, we ablated different choices of spatially dynamic self-attention by alternating window/ deformable/neighborhood attention. We removed positional encoding to eliminate its effect for now. Alternating window attention (WMSA) with both DMSA and NSMA led to a performance gain of $0.7\%$ (DMSA) and $0.4\%$ (NSMA), respectively. Alternating NMSA and DMSA resulted in a performance gain of $1.4\%$. Third, we brought back positional encoding, demonstrating that not all PEs can lead to a performance gain, e.g., adding APE and CPE even led to a performance drop. This is because these two PEs were not designed for irregularly sampled grids introduced by DMSA. Instead, the proposed MS-DePE improved performance by $0.8\%$. Notably, the proposed AgileFormer did not lead to a huge computational burden, with only a 1.1% increase in the number of parameters and a 15% rise in floating point operations, compared to SwinUNet.
3.4 Model scaling behavior
We compared the model scaling behavior of the proposed method with other ViT-UNets. As shown in Figure˜1(a), the proposed method demonstrated improved performance with growing model sizes for the Synapse multi-organ segmentation task, where different organs exhibit a lot of variability in shape and size. Specifically, the performance was improved by 2.15 in DSC (%) going from AgileFormer-T to AgileFormer-B, with parameters increase of approximately fourfold from 28.85M to 123.47M. We also observed that the methods with dynamic modules scaled better than those without dynamic modules, except for CoTr and SDAUT. We hypothesized this might be attributed to the fact that the main backbone of CoTr was still a CNN. Besides, deformable modules in both CoTr and SDAUT were only placed in the bottleneck.
4 Conclusion
In this paper, we proposed AgileFormer that systematically introduced spatially dynamic components (i.e., deformable patch embedding, spatially dynamic self-attention, and multi-scale deformable positional encoding) into a UNet for capturing spatially dynamic information for diverse target objects in the medical image segmentation. Extensive experiments demonstrated the effectiveness of the proposed method for a variety of medical image segmentation tasks. We aspire that the idea of systematically introducing spatially dynamic components to ViT-UNet will guide future designs on how to extract spatially dynamic representation for medical image segmentation with multiple targets.
Appendix
Figure 5: (a) CoTr is a hybrid model where Deformable DETER [37] was applied to the bottleneck. (b) SDAUT is a ViT-UNet originally proposed for medical image restoration. The encoder/decoder of SDAUT used window attention (WMSA); while DMSA was applied to the bottleneck. (c) MERIT captured dynamic feature representations using a multi-resolution and multi-stage approach, with the MaxViT [38] as the building block. (d) AgileFormer alternated NMSA and DMSA as the building block, adding novel deformable patch embedding and multi-scale deformable positional encoding.
Appendix A Comparison with other dynamic ViT-UNets
CoTr. Similar to its ancestor model [37], CoTr is a hybrid model where the main backbone (encoder/decoder) is still a CNN. Multi-scale deformable attention outlined in [37] serves as a bottleneck. As a result, it can only introduce limited deformability to capture spatially dynamic feature representation, as the main feature extractor does not have spatially dynamic components. In addition, CoTr does not take full advantage of ViT to capture long-range dependencies as its main backbone is still a CNN. We hypothesize these factors may limit its scalability (see evidence in Appendix C). Instead, the proposed method uses a ViT as the main backbone (encoder/decoder), where neighborhood attention and deformable attention [15] serve as the main building blocks, which better preserves the advantage of ViT while having the additional advantage of capturing spatially localized and dynamic features.
SDAUT. Although SDAUT also uses the deformable attention mechanism [15] as the proposed method, it is primarily proposed for image restoration tasks instead of a segmentation task. The optimal architecture of SDAUT uses a combination of window attention and deformable attention, where deformable attention is applied only to the third encoder and decoder. Instead, the deformable attention is applied at every layer (encoder/decoder) of the proposed method, which facilitates the extraction of spatially dynamic features. In addition, we further introduce deformable patch embedding to fundamentally change the rigid patch embedding. We also introduce a multi-scale deformable positional encoding to handle the irregularly sampled grid introduced by the deformable attention. We found that those two novel components are as important as spatially dynamic attention to enhance the segmentation performance.
MERIT. MERIT mainly takes advantage of multi-resolution information to capture spatially dynamic (mainly varying size) features, by combining the features extracted from inputs of different sizes. Instead, the proposed method handles both varying sizes as well as varying shapes by spatially dynamic attention. Whereas, we use a multi-scale deformable positional encoding to further capture multi-scale information.
Appendix B Implementation details on 3D experiments
We found that the comparison in nnFormer [20] was unfair. nnFormer and SwinUNETR take images with a size of $128\times 128\times 128$ as input; while the results reported by UNETR were achieved with inputs of a size of $96\times 96\times 96$ in our implementation. As a result, the input image size was set to $96\times 96\times 96$ for all methods. We also found that there was a huge performance gap between nnFormer and the other methods (i.e., TransUNet, TransBTS, UNETR, SwinUNETR, and CoTr), which is not consistent with the performance on other segmentation tasks. We conjectured that this performance gap might be attributed to different preprocessing, training, and evaluation procedures. Therefore, in our experiments, we retrained all the methods using the nnFormer codebase to standardize the comparison. For the proposed method, the batch size was set to 4, and the initial learning rate was set to $5\times 10^{-4}$. For the baseline methods, we followed the default hyperparameter settings in their original papers. We found out that our reimplmentation resulted in a higher performance of TransUNet, TransBTS, UNETR, SwinUNETR, and CoTr compared to results reported in nnFormer, even with a smaller size input.
[…]